OpenStudy (anonymous):

helppppp A survey of teenagers found the music preferences listed. If you pick a teenager at random, what is P(rock)? Country 12 Jazz 5 Rock 7 Classical 6 New Age 6 Other 4

OpenStudy (anonymous):

4/40 or 1/10

OpenStudy (anonymous):

What? those aren't even answer choices..

OpenStudy (anonymous):

Total number of teenagers = 12+5+7+6+6+4 = 40 Number of rock lovers = 7 So P(Rock/Teenagers) = 7/40.

OpenStudy (anonymous):

Thank youuuu(;

OpenStudy (anonymous):

ooops